OpenStudy (anonymous):

trying to verify.... using trig identities... 1/1+cosA + 1/1-cosA = 2 + 2cot^2a? I set it up this way for the r.h.s.... 1+ cos^2a/sin^2 + 1 + cos^2/sin^2a... is this correct?

OpenStudy (aum):

\[ \frac{1}{1+\cos A} + \frac{1}{1-\cos A} = \frac{1 -\cos A + 1 + \cos A}{1-\cos^2 A} = \\ \frac{2}{\sin^2 A} = 2\csc^2A = 2(1+\cot^2 A) = 2 + 2\cot^2 A \]

OpenStudy (aum):

\[\frac ab + \frac cd = \frac{ad + bc}{bd} \]

OpenStudy (aum):

\[(x+y)(x-y) = x^2 - y^2\]

OpenStudy (anonymous):

Wow, i thought the rhs was the more difficult side to prove. would it be possible if you could you set the right side equal to the left if you don't mind...thank you very much

OpenStudy (anonymous):

oh my gosh nevermind. im sorry im such an idiot. yes thank you i understand.

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

im sorry i had one more question, where does the 1/ 1+cos a + 1/1-cosa come from if it's 1-cos a + 1+cos a in the numerator?

OpenStudy (anonymous):

nevermind figured. it out thankyou again